# Fine-tuning an ECG Foundation Model to Predict Coronary CT Angiography Outcomes

### v1

@@ ABSTRACT
Coronary artery disease (CAD) remains a major global health burden. Accurate identification of the culprit vessel and assessment of stenosis severity are essential for guiding individualized therapy. Although coronary CT angiography (CCTA) is the first-line non-invasive modality for CAD diagnosis, its dependence on high-end equipment, radiation exposure, and strict patient cooperation limits large-scale use. With advances in artificial intelligence (AI) and the widespread availability of electrocardiography (ECG), AI-ECG offers a promising alternative for CAD screening. In this study, we developed an interpretable AI-ECG model to predict severe or complete stenosis of the four major coronary arteries on CCTA. On the internal validation set, the model’s AUCs for the right coronary artery (RCA), left main coronary artery (LM), left anterior descending artery (LAD), and left circumflex artery (LCX) were 0.794, 0.818, 0.744, and 0.755, respectively; on the external validation set, the AUCs reached 0.749, 0.971, 0.667, and 0.727, respectively. Performance remained stable in a clinically normal-ECG subset, indicating robustness beyond overt ECG abnormalities. Subgroup analyses across demographic and acquisition-time strata further confirmed model stability. Risk stratification based on vessel-specific incidence thresholds showed consistent separation on calibration and cumulative event curves. Interpretability analyses revealed distinct waveform differences between high- and low-risk groups, highlighting key electrophysiological regions contributing to model decisions and offering new insights into the ECG correlates of coronary stenosis.
KEYWORDS
Electrocardiogram, Coronary CT Angiography, Artificial Intelligence, Deep Learning, Coronary Heart Disease
@@ INTRODUCTION
Currently, coronary CT angiography (CCTA) is recommended as the first-line non-invasive modality for diagnosing CAD, providing a detailed assessment of coronary anatomy and stenosis severity, while serving as a crucial tool for pre-procedural planning and strategy guidance for percutaneous coronary intervention (PCI)5, 6. However, although CCTA is a non-invasive examination, the procedure requires sophisticated equipment and image post-processing techniques, and is highly dependent on patient heart rate control and respiratory coordination7. Furthermore, CCTA carries certain radiation exposure risks8 and may cause contrast-induced kidney injury9, limiting its application, especially in patients with renal insufficiency. Therefore, exploring a rapid, low-cost, non-invasive alternative that accurately reflects coronary artery lesions is of significant clinical importance for optimizing treatment processes, improving the efficiency of medical resource utilization, and enhancing patient outcomes.
@@
Electrocardiography (ECG), as a non-invasive, convenient, and real-time detection method, has unique advantages in the early identification and risk stratification of cardiovascular diseases. Changes in ECG waveforms, such as ST segment shift and T wave inversion, have been widely used in the preliminary diagnosis of acute coronary syndrome (ACS)10, 11, 12. Meanwhile, with the rapid development of artificial intelligence (AI), especially deep learning technology, AI is being widely used in clinical research 13, 14, 15, 16, 17, 18. Utilizing AI to perform deep feature mining of ECG signals to predict the pathogenesis of cardiovascular diseases has become a research hotspot.19, 20, 21, 22, 23. However, existing AI-ECG studies on CAD-related coronary artery disease using CCTA mostly focus on whether there is screening for obstructive coronary artery disease. For example, Choi et al.’s study mainly targeted the detection of obstructive coronary artery disease24, while Park et al.’s study focused on predicting obstructive lesions in patients with stable angina25. Although some studies have attempted to use ECG and AI models to locate myocardial infarction (MI) sites, their research subjects are mostly scar tissue from previous myocardial infarctions, lacking systematic research on the localization of the responsible vessel during acute coronary syndrome and the identification of coronary artery occlusion on normal ECG 26.
To address the aforementioned shortcomings, this study integrated clinical data from multiple centers, including Peking University People’s Hospital and the Second Hospital of Tianjin Medical University. Employing a transfer learning strategy, we fine-tuned ECGFounder27, 28, the world’s largest ECG foundation model trained for predicting future new-onset MI. We constructed a predictive model that can directly infer the diagnostic results of severe stenosis and complete occlusion shown by CCTA from ECG signals, and further implemented a stratified assessment of the risk of future new-onset stenosis in major coronary arteries. Simultaneously, we systematically evaluated the model’s performance in a population with normal ECG, providing a new technical approach for the early screening and risk identification of occult CAD. Furthermore, this study deeply analyzed the changing characteristics of ECG waveforms in different risk populations, revealing potential signal patterns related to coronary artery stenosis from an electrophysiological perspective, providing a new perspective for understanding its pathogenesis, and further expanding the application boundaries of ECG in the precise diagnosis of coronary artery stenosis.
Overall, the principal innovations and contributions of this study are summarized as follows.
  1. 1.
    We developed an interpretable AI-ECG model capable of directly predicting CCTA-defined severe coronary stenosis and complete occlusion from ECG signals. The predictive performance and probability calibration of the model were systematically evaluated using receiver operating characteristic (ROC) curves, box plots of predicted probabilities across different stenosis severities with statistical testing, and calibration curves.
  2. 2.
    We further investigated the model’s ability to identify CCTA-defined severe stenosis and complete occlusion in individuals with apparently normal ECGs, demonstrating its potential value for the detection of concealed CAD.
  3. 3.
    By integrating model-predicted probabilities with observed incidence rates, we stratified individuals into high-risk and low-risk groups and constructed cumulative incidence curves, enabling dynamic risk stratification for future coronary events.
  4. 4.
    Through explainable analysis, we comprehensively compared ECG waveform characteristics between model-defined high-risk and low-risk populations, providing novel insights into the electrocardiographic manifestations associated with coronary artery stenosis and offering a new perspective for clinical interpretation.
@@ RESULTS
This study used 2,595 CCTA examinations and 4,620 paired ECG data from 2,323 patients at Peking University People’s Hospital, and 135 CCTA examinations and 228 paired ECG data from 135 patients at the Second Hospital of Tianjin Medical University. We used the data from Peking University People’s Hospital as the internal dataset and the data from the Second Hospital of Tianjin Medical University as the external dataset to validate the model’s generalization ability on multicenter data.
Table 1 shows the baseline characteristics of the internal and external data used in the study.
Table 1: Baseline Characteristics of Internal and External Cohorts
Characteristic Peking University People’s Hospital (Internal) The Second Hospital of Tianjin Medical University (External)
No. of Patients 2323 135
@@
      Mean $\pm$ SD $63.7\pm 10.1$ $62.6\pm 14.2$
      Median (Range) 64.0(21.092.0)64.0\ (21.0-92.0) 65.0(28.093.0)65.0\ (28.0-93.0)
Sex (n (%))
      Male $1042\ (55.0\%)$ $80\ (61.1\%)$
      Female $853\ (45.0\%)$ $51\ (38.9\%)$
@@
We used patient ID as the grouping criterion and employed a five-fold stratified grouping cross-validation strategy to train and validate the AI-ECG model while ensuring a balanced distribution of multi-labels across each fold. To systematically evaluate the model’s generalization performance on a multi-center dataset and its predictive ability in patients with clinically diagnosed normal ECGs, this study selected the best-performing fold result from the internal validation set, the overall result from the external validation set, and the prediction result of the best-performing fold for patients with ECG diagnoses of normal. Three ROC curves were plotted, as shown in Figure 1.
The results show that, in the internal validation set, the predicted AUCs for the right coronary artery (RCA), left main coronary artery (LM), left anterior descending artery (LAD), and left circumflex artery (LCX) are 0.794, 0.818, 0.744, and 0.755, respectively; while in the external validation set, the corresponding AUCs are 0.749, 0.971, 0.667, and 0.727. These results demonstrate that our AI-ECG model exhibits stable and superior discriminative performance in identifying severe stenosis and complete occlusion of the four major coronary arteries. The external validation results are generally consistent with the internal validation results, further confirming the model’s good generalization ability across different central datasets.
Furthermore, we evaluated the model’s predictive ability in the internal validation set for individuals with clinically normal ECGs. The results showed that in this subgroup, the predictive AUCs for RCA, LM, LAD, and LCX were 0.793, 0.896, 0.707, and 0.765, respectively. These results are largely consistent with the model’s performance in the full population within the internal validation set, indicating that even in individuals with superficially normal ECGs, the AI-ECG model can effectively identify the potential risk of severe coronary artery stenosis, highlighting its potential application value in screening for occult CAD.
@@
(a) Internal Validation
(b) External Validation
(c) Normal ECG Prediction
Figure 1: ROC curves for model performance evaluation. (a) ROC curve validated using internal data from Peking University People’s Hospital. (b) External validation ROC curve of the Second Hospital of Tianjin Medical University. (c) Predicted ROC curves of normal electrocardiograms from internal data.
The study further conducted subgroup analyses in the internal validation set to assess the model’s stability across different population characteristics and time conditions. After stratifying the population according to age (<65years<65years vs. $\geq 65years$), sex, and the time interval between ECG and CCTA examinations ($\leq 3hours$ vs. >3hours>3hours), the model’s discriminative performance remained consistent across all subgroups.
As shown in Figure 2, the AUC levels of the model remained generally stable across subgroups, with no significant performance decline.
Specifically, the model exhibited a relatively higher AUC in the subgroup where the interval between ECG and CCTA examinations was greater than 3 hours; it also demonstrated superior discriminative performance in the younger group (under 65 years of age). Between the male and female subgroups, the model’s predictive performance was largely consistent, with no significant differences.
These results further demonstrate that the model possesses good stability and robustness under different population characteristics and testing conditions, further supporting its potential application value in diverse clinical scenarios.
Figure 2: Subgroup analyses. Subgroup analysis results demonstrate that the model’s predictive performance remains robust across different age groups, genders, and time intervals.
@@
Spearman rank correlation analysis confirmed a statistically significant positive correlation between the predicted probability and the degree of stenosis in all coronary arteries ($p<0.001$), with the left anterior descending artery (LAD) showing the strongest correlation ($\rho=0.435$). Notably, the absolute predicted probability of the left main coronary artery (LM) (range: 0.0–0.2) was significantly lower than that of other vessels (range: 0.0–0.8), reflecting the low prevalence of LM lesions in the training population. However, the model still maintained significant discriminative power for LM ($\rho=0.307$).
@@
Figure 4: Calibration curve of model prediction results. The calibration curves demonstrate the model’s reliability in predicting each vessel, and all four coronary arteries exhibited low Brier scores.
@@
The calibration curves show that the model demonstrates good reliability in predicting all four coronary arteries. The Bridge scores for RCA, LM, LAD, and LCX are 0.112, 0.030, 0.166, and 0.106, respectively. Notably, for severe and complete occlusion of the LM, the model did not output a very high prediction probability due to the extremely low incidence of this lesion in the general population (approximately 0.1%), which is consistent with epidemiological distribution.
@@
To clearly demonstrate its performance, we magnified the LM subplot locally (between 0 and 0.3). The results show that although the predicted probabilities are concentrated in the low range, the high-risk group still exhibits a clear upward trend in risk, and the Brier Score (0.030) is extremely low, proving that the model has extremely high accuracy in excluding the vast majority of healthy people (High Negative Predictive Value), and does not produce a large number of false positives due to blindly pursuing high probabilities.
To further validate the model’s decision-making ability in clinical risk stratification, we plotted the cumulative event occurrence curves based on the model’s predicted risk stratification, as shown in Figure 5. Using the ECG monitoring time as the follow-up starting point, we defined the occurrence of severe stenosis and complete occlusion as the event endpoints, with a follow-up period of one year. Individuals who did not develop severe stenosis or complete occlusion during the one-year follow-up period were defined as having no events. By comparing the cumulative event occurrence curves of high-risk and low-risk populations, we assessed the model’s ability to distinguish future coronary events, thereby validating its clinical risk stratification value over time.
Figure 5: Cumulative event occurrence curve based on model-predicted risk stratification. The cumulative event occurrence curve shows that the stratification results for each coronary artery exhibit a good separation trend and significant statistical differences.
As shown in Figure 5, the cumulative event occurrence curves of the high-risk and low-risk groups in the four major coronary arteries exhibit a clear and stable separation trend, demonstrating the model’s good ability to distinguish between different risk levels. The differences between the two groups were statistically significant (log-rank test, all P < 0.001), further validating the model’s discriminative power in coronary event risk stratification and its potential clinical application value.
Interpretability Analysis Results
To further reveal the key ECG regions that the model focuses on in risk assessment and stratification, we conducted an interpretability analysis based on the waveform level. After dividing the population into high-risk and low-risk groups according to the model’s predicted probabilities, we extracted normalized individual heartbeat segments from each patient’s raw ECG signals and aligned and standardized them. Subsequently, within each risk group, we calculated the mean morphology and standard deviation of the heartbeat waveforms in each lead and presented them in a multi-lead visualization.
@@
As shown in Figure 6, by comparing the differences in waveform morphology between high-risk and low-risk groups, the electrocardiographic feature patterns on which the model distinguishes different coronary artery risk states are intuitively revealed, providing interpretable evidence for understanding its discrimination mechanism and potential electrophysiological basis.
@@
Figure 6: Interpretability analysis results. The results reveal the key ECG regions that the model focuses on during prediction and stratification.

### v2

@@ Abstract
Coronary artery disease (CAD) remains a major global public health burden, yet scalable tools for risk screening are limited. Although coronary computed tomography angiography (CCTA) is a first-line non-invasive diagnostic modality, its widespread use is constrained by resource requirements and radiation exposure. Artificial intelligence–enabled electrocardiography (AI-ECG) may provide a complementary approach for CAD risk stratification. We developed and validated an AI-ECG model using CCTA as the reference standard to estimate severe ($\geq 70\%$) or complete ($\geq 99\%$) stenosis in the four major coronary arteries. In internal validation, the model achieved area under the receiver operating characteristic curve (AUC) values of 0.706–0.744 across vessels and demonstrated consistent performance in external validation (AUCs: 0.673–0.714). Discrimination remained stable among individuals with clinically normal ECGs and across demographic and clinical subgroups. In a dedicated clinical cohort with longitudinal follow-up, vessel-specific risk stratification based on model-predicted probabilities yielded distinct separation between high-risk and low-risk groups in time-to-event analyses using Kaplan–Meier curves, while decision curve analysis suggested potential clinical utility as an adjunctive screening tool. Explainable analyses highlighted waveform patterns associated with elevated predicted risk. These findings support the feasibility of AI-ECG for complementary CAD risk screening and warrant prospective evaluation.
@@ Introduction
Currently, coronary CT angiography (CCTA) is recommended as the first-line non-invasive modality for diagnosing CAD, providing a detailed assessment of coronary anatomy and stenosis severity, while serving as a crucial tool for pre-procedural planning and strategy guidance for percutaneous coronary intervention (PCI)5, 6. Nevertheless, the widespread application of CCTA is limited by its dependence on advanced imaging infrastructure and post-processing techniques, strict requirements for heart rate control and respiratory coordination, radiation exposure, and the risk of contrast-induced nephropathy, particularly in patients with impaired renal function 7, 8, 9. These limitations underscore the need for complementary, low-cost, and non-invasive approaches that can assist in CAD risk assessment and triage.
@@
Electrocardiography (ECG), as a non-invasive, convenient, and real-time detection method, has unique advantages in the early identification and risk stratification of cardiovascular diseases. Changes in ECG waveforms, such as ST segment shift and T wave inversion, have been widely used in the preliminary diagnosis of acute coronary syndrome (ACS)10, 11, 12. Meanwhile, with the rapid development of artificial intelligence (AI), especially deep learning technology, AI is being widely used in clinical research 13, 14, 15, 16, 17, 18. Utilizing AI to perform deep feature mining of ECG signals to predict the pathogenesis of cardiovascular diseases has become a research hotspot 19, 20, 21, 22, 23. However, existing AI-ECG studies on CAD-related coronary artery disease using CCTA mostly focus on whether there is screening for obstructive coronary artery disease. For example, Choi et al.’s study mainly targeted the detection of obstructive coronary artery disease24, while Park et al.’s study focused on predicting obstructive lesions in patients with stable angina25. Although some studies have attempted to use ECG and AI models to locate myocardial infarction (MI) sites, their research subjects are mostly scar tissue from previous MI, lacking systematic research on the localization of the responsible vessel during acute coronary syndrome and the identification of coronary artery occlusion on normal ECG 26.
To overcome these limitations, we conducted a multicenter study that included clinical data from Peking University People’s Hospital and the Second Hospital of Tianjin Medical University, as well as a dedicated longitudinal follow-up cohort. We employed a transfer learning framework to fine-tune ECGFounder, a large-scale pretrained electrocardiogram foundation model 27, 28 which was adapted by our group on an independent dataset for MI–related prediction tasks, to develop an AI-ECG model capable of assessing severe stenosis and total occlusion of the four major coronary arteries as defined by coronary CT angiography (CCTA). Furthermore, we implemented vessel-specific risk stratification based on model-predicted probabilities and evaluated performance in individuals with clinically normal ECGs. To further explore the electrophysiological relevance underlying the model’s predictions, we performed explainable analyses comparing waveform characteristics between high-risk and low-risk groups.
Overall, this study investigated the development and evaluation of an explainable artificial intelligence electrocardiogram (AI-ECG) model for the direct prediction of coronary computed tomography angiography (CCTA)-defined severe coronary artery stenosis and total occlusion from standard ECG signals. Model performance was systematically assessed using receiver operating characteristic analyses, comparisons of predicted probability distributions across different stenosis severities, and calibration analyses.
The model was further evaluated in individuals with apparently normal electrocardiograms to examine its behavior in populations without overt electrocardiographic abnormalities. In addition, model-predicted probabilities were integrated with epidemiological incidence data to perform vessel-specific risk stratification and to assess cumulative myocardial infarction risk over time using Kaplan–Meier analyses. Sensitivity analyses across a range of decision thresholds were conducted to explore operating points relevant to opportunistic screening scenarios. Finally, explainability analyses were performed to characterize differences in ECG waveform patterns between model-defined high-risk and low-risk groups, with the aim of supporting clinical interpretation.
@@ Results
This study included multicenter clinical data from Peking University People’s Hospital and the Second Hospital of Tianjin Medical University. Peking University People’s Hospital included 2323 patients, comprising 2595 coronary CT angiography (CCTA) examinations and 4620 paired electrocardiogram (ECG) records; the Second Hospital of Tianjin Medical University included 2477 patients, comprising 2477 CCTA examinations and 2477 paired ECG records. Furthermore, we constructed a longitudinal follow-up cohort from the Second Hospital of Tianjin Medical University, including 400 patients, to evaluate the model’s predictive ability in future risk stratification.
Data from Peking University People’s Hospital was used as the internal dataset, and data from the Second Hospital of Tianjin Medical University was used as the external validation dataset to systematically evaluate the model’s generalization performance in a multicenter scenario; the longitudinal follow-up cohort was used to validate the model’s ability to stratify the risk of future adverse events (MI).
Table 1 shows the baseline characteristics of the data used in the study.
Table 1: Baseline Characteristics of Internal and External Datasets and Cohort
Characteristic Peking University People’s Hospital (Internal) The Second Hospital of Tianjin Medical University (External) Cohort of the Second Hospital of Tianjin Medical University
($n=4,620$) ($n=2,477$) ($n=400$)
Demographics
Sex
     Male 1,952 (42.3%) 1,283 (51.8%) 183 (45.8%)
     Female 1,922 (41.6%) 1,194 (48.2%) 217 (54.2%)
     Missing 746 (16.1%) 0 (0.0%) 0 (0.0%)
@@
     Mean $\pm$ SD 64.5 $\pm$ 10.4 63.1 $\pm$ 11.7 63.14 $\pm$ 10.43
     <65<65 1,886 (40.8%) 1,223 (49.4%) 199 (49.8%)
     $\geq 65$ 1,988 (43.0%) 1,252 (50.5%) 201 (50.2%)
     Missing 746 (16.1%) 2 (0.1%) 0 (0.0%)
Medical History
Diabetes
     Yes 1,433 (31.0%) 889 (35.9%) 97 (24.3%)
     No 3,187 (69.0%) 1,588 (64.1%) 303 (75.7%)
Hypertension
     Yes 2,411 (52.2%) 1,256 (50.7%) 171 (42.8%)
     No 2,209 (47.8%) 1,221 (49.3%) 229 (57.2%)
Prior myocardial infarction
     Yes 193 (4.2%) 49 (2.0%) 16 (4.0%)
     No 4,427 (95.8%) 2,428 (98.0%) 384 (96.0%)
Coronary Artery Stenosis
RCA
     Severe stenosis & occlusion 795 (17.2%) 635 (25.6%) 64 (16.0%)
     Non-severe or normal 3,825 (82.8%) 1,842 (74.4%) 336 (84.0%)
LM
     Severe stenosis & occlusion 235 (5.1%) 52 (2.1%) 4 (1.0%)
     Non-severe or normal 4,385 (94.9%) 2,425 (97.9%) 396 (99.0%)
LAD
     Severe stenosis & occlusion 1,373 (29.7%) 769 (31.0%) 83 (20.8%)
     Non-severe or normal 3,247 (70.3%) 1,708 (69.0%) 317 (79.2%)
LCX
     Severe stenosis & occlusion 795 (17.2%) 570 (23.0%) 55 (13.8%)
     Non-severe or normal 3,825 (82.8%) 1,907 (77.0%) 345 (86.2%)
  • Data are presented as n (%) or mean $\pm$ SD.
  • Abbreviations: SD, standard deviation; RCA, right coronary artery; LM, left main coronary artery; LAD, left anterior descending artery; LCX, left circumflex artery.
@@
We employed a five-fold stratified cross-validation strategy with patient identifiers used as the grouping variable, ensuring that all ECG records from the same individual were assigned to a single fold and that the distribution of multi-label coronary artery outcomes was balanced across folds. To comprehensively evaluate model generalizability across multicenter datasets and its performance among individuals with clinically normal ECGs, we reported micro-averaged area under the receiver operating characteristic curve (micro-AUC) for the internal dataset, as well as performance on the external validation dataset using an ensemble of models derived from the five cross-validation folds. Model performance was additionally assessed in the subgroup of patients with clinically interpreted normal ECGs. Receiver operating characteristic (ROC) curves corresponding to these three evaluation settings are presented in Figure 1.
The results show that, in the internal validation set, the predicted AUCs for the right coronary artery (RCA), left main coronary artery (LM), left anterior descending artery (LAD), and left circumflex artery (LCX) are 0.744, 0.706, 0.716, and 0.736, respectively; while in the external validation set, the corresponding AUCs are 0.714, 0.713, 0.700, and 0.673. These results demonstrate that our AI-ECG model exhibits stable and superior discriminative performance in identifying severe stenosis and complete occlusion of the four major coronary arteries. The external validation results are generally consistent with the internal validation results, further confirming the model’s good generalization ability across different central datasets.
Furthermore, we evaluated the model’s predictive ability in the internal validation set for individuals with clinically normal ECGs. The results showed that in this subgroup, the predictive AUCs for RCA, LM, LAD, and LCX were 0.693, 0.659, 0.673, and 0.716, respectively. These results are largely consistent with the model’s performance in the full population within the internal validation set, indicating that even in individuals with superficially normal ECGs, the AI-ECG model can effectively identify the potential risk of severe coronary artery stenosis, highlighting its potential application value in screening for occult CAD.
@@
Figure 1: Model discrimination performance assessed by receiver operating characteristic curves. a) Internal validation using data from Peking University People’s Hospital. b) External validation using data from the Second Hospital of Tianjin Medical University. c) Performance evaluation in individuals with clinically normal electrocardiograms from the internal dataset. Area under the curve (AUC) values are reported with 95% confidence intervals (CI).
In addition, to account for potential differences between examination-level and patient-level evaluation, ROC analyses were performed and reported at both the CCTA examination level and the patient level, with the corresponding curves shown in Figure 2.
Figure 2: ROC analyses at the CCTA examination and patient levels. ROC curves are shown for model performance evaluated at both the CCTA examination level and the patient level, illustrating consistency of discrimination across different evaluation granularities.
Subgroup analyses were further performed to evaluate model stability across population characteristics and acquisition timing. After stratification by age (<65years<65years vs. $\geq 65years$), sex, coronary artery stenosis, the interval between ECG and CCTA acquisition ($\leq 3hours$ vs. >3hours>3hours), and the medical history, the model demonstrated largely comparable discriminative performance across subgroups, without evidence of substantial performance degradation.
As shown in Figure 3, model performance remained largely consistent across prespecified subgroups, with no substantial decline in area under the receiver operating characteristic curve (AUC) observed, except among individuals with a prior history of MI.
Specifically, higher AUC values were observed in the subgroup with an ECG–CCTA acquisition interval exceeding 3 hours, as well as among younger individuals (<65 years) and those with a history of hypertension. Model discrimination was comparable between male and female participants and did not differ materially according to diabetes status.
Collectively, these findings suggest that the proposed model demonstrates stable discriminative performance across diverse population characteristics and testing conditions, supporting its potential applicability in heterogeneous clinical settings.
Subgroup analyses stratified by the severity of coronary artery stenosis are provided in Table A1.
Figure 3: Subgroup analyses of model performance. Forest plots show the AUC with 95% CI across clinically relevant subgroups, including age, sex, time from symptom onset, and medical history.
@@
Spearman rank correlation analysis confirmed a statistically significant positive correlation between the predicted probability and the degree of stenosis in all coronary arteries ($p<0.001$), with the RCA showing the strongest correlation ($\rho=0.371$). Notably, the absolute predicted probability of the left main coronary artery (LM) (range: 0.0–0.4) was significantly lower than that of other vessels, reflecting the low prevalence of LM lesions in the training population. However, the model still maintained significant discriminative power for LM ($\rho=0.240$).
@@
Figure 5: Calibration curve of model prediction results. Calibration plots illustrate the agreement between predicted and observed risks for each coronary artery. Red circles indicate individuals classified as the high-risk group, whereas blue circles indicate individuals classified as the low-risk group. All four coronary arteries demonstrated good calibration performance with low Brier scores.
@@
The calibration curves show that the model demonstrates good reliability in predicting all four coronary arteries. The Brier scores for RCA, LM, LAD, and LCX are 0.130, 0.047, 0.190, and 0.131, respectively. Notably, for severe and complete occlusion of the LM, the model did not output a very high prediction probability due to the extremely low incidence of this lesion in the general population (approximately 0.1%), which is consistent with epidemiological distribution.
@@
To clearly demonstrate its performance, we magnified the LM subplot locally (between 0 and 0.35). The results show that although the predicted probabilities are concentrated in the low range, the high-risk group still exhibits a clear upward trend in risk, and the Brier Score (0.047) is extremely low, proving that the model has extremely high accuracy in excluding the vast majority of healthy people (High Negative Predictive Value), and does not produce a large number of false positives due to blindly pursuing high probabilities. The positive and negative rates of the experimental results are shown in Table A2.
To further evaluate the model’s potential value in clinical risk stratification, we grouped patients based on model-predicted probabilities. Furthermore, we conducted a longitudinal comparison of the risk of myocardial infarction at different risk levels, as shown in Figure 6. The time of ECG acquisition was used as the start of follow-up, and an event was defined as a myocardial infarction occurring during follow-up.
The analysis focused on incident events occurring within a predefined follow-up window of 700 days. Through Kaplan-Meier analysis, we compared the cumulative risk of myocardial infarction between the high-risk and low-risk groups during follow-up, thereby assessing the model’s ability to distinguish changes in future myocardial infarction risk over time.
Figure 6: Kaplan–Meier curves stratified by model-predicted risk across major coronary arteries. Kaplan–Meier plots show the cumulative incidence of myocardial infarction over the follow-up period for individuals stratified into high-risk and low-risk groups based on model-predicted probabilities across the four major coronary arteries. Red solid lines represent the high-risk group, whereas blue solid lines represent the low-risk group.
As shown in Figure 6, Kaplan-Meier survival curves in all four major coronary arteries revealed a significant separation between high-risk and low-risk groups during follow-up. The cumulative incidence of myocardial infarction gradually increased over time, indicating a graded risk pattern in the risk stratification defined by the model. Log-rank tests showed statistically significant differences between risk groups in RCA (p < 0.001), LAD (p = 0.029), and LCX (p = 0.002). A trend of risk separation was also observed in the LM curve. In summary, these results demonstrate that the model can capture meaningful longitudinal risk stratification signals in major coronary arteries.
To evaluate the potential clinical utility of the proposed risk stratification strategy, we further conducted decision curve analysis (DCA) to quantify the net benefits of model-guided decision-making at different clinically relevant risk thresholds. As shown in Figure 7, in the four major coronary arteries, the DCA curves corresponding to the model are generally above the treat-all and treat-none strategies, indicating that model-based risk stratification can achieve higher net clinical benefits within the corresponding threshold ranges, supporting its potential application value as an auxiliary decision-making tool.
Figure 7: Decision curve analysis for model-guided coronary risk stratification. Decision curve analysis (DCA) illustrates the net clinical benefit of model-guided decision-making across a range of clinically relevant risk thresholds for the four major coronary arteries.
To further evaluate robustness, we examined model sensitivity and specificity across multiple alternative risk thresholds in addition to those used in the primary analysis. Performance remained consistent across a reasonable range of thresholds, indicating potential operating points for opportunistic screening (Figure A1 and Table A3).
Explainability Analysis Results
To further elucidate the electrocardiographic regions contributing to model-based risk assessment and stratification, we performed an explainable analysis at the waveform level. Following stratification of the population into high-risk and low-risk groups based on model-predicted probabilities, individual heartbeat segments were extracted from raw ECG recordings, normalized, and temporally aligned. R-wave locations were identified on lead II using a peak detection algorithm, with predefined amplitude thresholds and a minimum RR interval to ensure robust detection of dominant QRS complexes. Using the detected R-wave as a reference point, standardized heartbeat waveforms were obtained for each individual. Within each risk group, the mean waveform morphology and corresponding standard deviation were computed for each lead and visualized in a multi-lead format to facilitate comparison between risk strata.
@@
As shown in Figure 8, by comparing the differences in waveform morphology between high-risk and low-risk groups, the electrocardiographic feature patterns on which the model distinguishes different coronary artery risk states are intuitively revealed, providing explainable evidence for understanding its discrimination mechanism and potential electrophysiological basis.
@@
Figure 8: Explainable analysis results. The results reveal the key ECG regions that the model focuses on during prediction and stratification. Red lines indicate individuals classified as the high-risk group, whereas blue lines indicate individuals classified as the low-risk group.